QUESTION 19 21 Evaluate the finite geometric series : 2 2(3/5)-1 n=1 21 -50 14.9 4.99
Answer #1

21 껴 &= a (gni (0-0) S-au S here 0=27=(3) m=21 $ = 2 (C3)?-1) 5 $ = 4.99 please Upvote to: appreciołe my efforts